# Rethinking Latent Redundancy in Behavior Cloning: An Information Bottleneck Approach for Robot Manipulation

### v3

@@ 1 Introduction
Behavior Cloning (BC), one of the simplest and most widely used methods in Imitation Learning (IL), learns a mapping from states to actions by training on state-action pairs from expert demonstrations. BC has been widely studied in autonomous driving (Bain & Sammut, 1995; Torabi et al., 2018), robotics control (Argall et al., 2009) and game AI (Pearce & Zhu, 2022). In robot manipulation, BC has become a foundational approach, enabling robots to replicate expert actions based on sensory inputs such as images or proprioception information like gripper states. To enhance the generalization of robots, most BC methods focus on incorporating large datasets of human or manipulation videos (Jang et al., 2022; Karamcheti et al., 2023; Brohan et al., 2023; Cheang et al., 2024; Saxena et al., 2025), or integrating additional text and visual information (Jia et al., 2024; Wen et al., 2024; Hu et al., 2024). While these methods have made significant progress in improving generalization by leveraging more diverse information, they often neglect a critical aspect: whether the learned representations contain significant redundant information.
@@ 2.1 Behavior Cloning in Robot Manipulation
Behavioral Cloning (BC), first introduced by (Pomerleau, 1991), is a well-known Imitation Learning (IL) algorithm that learns a policy by directly minimizing the discrepancy between the agent’s actions and those of the expert in the demonstration data. To learn more generalizable representations, one class of visual representation learning methods pre-trains on large video datasets of robotics or humans, enabling rapid application of the pre-trained encoder to downstream robotic tasks. Notable examples include VC-1 (Majumdar et al., 2023), R3M (Nair et al., 2023), and Voltron (Karamcheti et al., 2023) . Meanwhile, another line of research focuses on training on even more extensive and diverse datasets with larger models, such as Internet-scale visual question answering and robot trajectory data (Brohan et al., 2023), as well as a vast collection of Internet videos (Cheang et al., 2024). Additionally, some methods further enhance generalization by incorporating additional sources of information. These include inferring textual descriptions based on the robot’s current state (Zawalski et al., 2024), leveraging visual trajectories (Wen et al., 2024) and generated images (Tian et al., 2025), and integrating 3D visual information (Goyal et al., 2023). However, these methods have not deeply analyzed the redundancy in learned latent representations, and most also lack a solid theoretical foundation. Thus we extend the Information Bottleneck (IB) principle to BC, addressing this fundamental gap.
@@ 4.1 Model Architecture
As illustrated in Figure 2 (b), spatial fusion involves extracting spatial features from data at a given time step or concatenating features across multiple time steps along the feature dimensions. This approach does not explicitly differentiate between time steps but instead processes the aggregated features as a whole, emphasizing the modeling of inter-feature relationships. The spatial fusion module can be implemented using Multi-Layer Perceptrons (MLPs), Convolutional Neural Networks (CNNs), Spatial Transformers, or even simple concatenation operations. These methods are primarily designed to learn highly generalizable visual encoders by leveraging large-scale human video datasets. The pretrained encoders are then fine-tuned for downstream robotic tasks (Nair et al., 2023; Karamcheti et al., 2023; Majumdar et al., 2023; Zeng et al., 2024).
@@
The latent representation $Z$, which integrates both spatial and temporal information, is then passed through a policy head to generate actions. Existing policy heads primarily focus on using MLPs, Gaussian Mixture Model (GMM), and diffusion-based policy (DP) heads (Chi et al., 2023; Reuss et al., 2024). For simplicity and clearer empirical demonstration, we use an MLP as the policy head.
@@ 4.2 Behavior Cloning with Information Bottleneck
where $z_{t}=F(x_{t})$ and $F(\cdot)$ denotes the fusion module.
@@ 5.1 Embodied Evaluation
Implementation. In CortexBench, for four partial fine-tuning methods, we train for 100 epochs on each task using the Adam optimizer with a learning rate of 1e-3, a batch size of 512, and weight decay of 1e-4, with learning rate decay applied using a cosine annealing schedule. For two full fine-tuning methods, we train for 50 epochs with a learning rate of 1e-4 and a batch size of 256. In LIBERO, we train for 50 epochs using the AdamW optimizer with a learning rate of 1e-4 and a batch size of 64, decayed using a cosine annealing schedule. In real-world evaluation, we train VC-1 for 200 epochs using the Adam optimizer with a learning rate of 1e-3 and a batch size of 512. CogAct is trained for 8k steps with the AdamW optimizer, using a learning rate of 2e-5 and a batch size of 128. For BC+IB methods, the model used in MINE consists of a two-layer MLP, with a learning rate of 1e-5. The Lagrange multiplier in Equation 6 ranges from 1e-4 to 5e-3 in this work.
@@ 5.3 Performance on LIBERO
Finding 8: IB is particularly effective for tasks requiring diverse feature extraction, such as distinguishing distinct task objectives or differentiating between various objects, as in LIBERO-Goal and LIBERO-Object. By filtering out irrelevant information, IB facilitates better generalization and more compact representations. However, the impact is less pronounced in spatial tasks such as LIBERO-Spatial, which rely heavily on structural information that can be disrupted by excessive compression. For long-horizon tasks, the main performance bottleneck lies in the lightweight baseline models, whose limited capacity restricts their effectiveness on more complex tasks such as LIBERO-Long. This limitation is further illustrated and solved in Appendix Table 5.
@@ 5.5 More Analysis
Effect of the Lagrange multiplier $\beta$ of Equation 6. This experiment evaluates how incorporating IB enhances performance. Since the MINE model’s parameters are fixed, the key difference between BC+IB and BC lies in the parameter $\beta$, which balances compression and predictive power. For LIBERO experiments, $\beta$ is explored within 1e-4, 1e-3, 5e-3, 1e-2. Finding 9: As shown in Figure 5, IB improves performance within a specific $\beta$ range, with a peak observed at an undetermined value. However, across all experiments, $\beta$ around 1e-4 consistently yields stable improvements. The selected $\beta$ values for each experiment are detailed in C.1.
@@ 6 Limitations and Discussion
While our work provides extensive experimental validation of the effectiveness of IB and the necessity of input redundancy reduction in robotics representation learning, several limitations remain. First, we do not comprehensively investigate the scalability of IB in robotic models. Most of our experiments are conducted on relatively lightweight architectures, with only the real-world experiments utilizing CogAct. We have not systematically evaluated large models such as vision-language-action architectures, due to the high computational and time costs involved. Second, alternative policy heads such as transformer-based designs, and models like RT-2 (Brohan et al., 2023) and OpenVLA  (Kim et al., 2024), which remove the explicit policy head and instead treat actions as text tokens, have not been explored. We leave these directions for future work. Third, although we evaluate our method on various benchmarks, its robustness to domain shifts, including variations in environment and task settings, remains underexplored. We hope our work will inspire future research and contribute to the ongoing development and refinement of these methods.
7 Conclusions
In this study, we investigated the redundancy in representations for Behavior Cloning in robot manipulation and introduced the Information Bottleneck principle to mitigate this issue. By incorporating IB, we aimed to filter out redundant information in latent representations while preserving task-relevant features. Extensive experiments across various representation learning methods on CortexBench and LIBERO revealed insightful findings and demonstrated that IB consistently improves performance across diverse tasks and architectures. We hope our work will inspire further integration of information-theoretic principles into robotics and foster deeper theoretical analysis in this domain.
@@ Acknowledgements
This work was supported by the National Natural Science Foundation of China with grant numbers (U21A20485 and 62436005), and the Fundamental Research Funds for Xi’an Jiaotong University under Grants xzy022024012.
@@
This paper presents work whose goal is to advance the field of Machine Learning in Robotics. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ B.2.2 Baselines in LEBERO
Lagrange multiplier $\beta$ [1e-4, 5e-3]
@@ B.3.2 LIBERO
For BC+IB, all training and model parameters remain identical to those of BC, except for the IB-specific parameters. The details of IB-related parameters are provided in Table 5, while the specific values of the Lagrange multiplier are thoroughly discussed in Section C.1.
@@ B.4.2 LIBERO
Table 6: Task-wise Performance on Adroit and Trifinger of CortexBench. We evaluated 4 tasks of 2 benchmarks using 3 random seeds and reported the average success rate (sr) and Lagrange multiplier $\beta$. The best performance is bolded.
@@
ResNet sr 65.33 66.67 66.00 87.12 56.06 71.59
ResNet+IB sr 69.33 74.67 72.00 87.14 57.45 72.30
$\beta$ 5e-3 1e-4 -- 5e-3 5e-3 --
ViT sr 61.33 9.33 35.33 78.77 32.37 55.57
ViT+IB sr 64.00 10.67 37.33 77.83 34.04 55.93
$\beta$ 5e-3 5e-3 -- 5e-3 5e-3 --
R3M sr 45.33 5.33 25.33 74.29 45.45 59.87
R3M+IB sr 52.00 2.67 27.33 75.04 46.23 60.63
$\beta$ 5e-3 5e-3 -- 5e-3 5e-3 --
Voltron sr 32.00 5.33 18.67 86.37 62.04 74.21
Voltron+IB sr 38.67 4.00 21.33 86.62 63.61 75.12
$\beta$ 5e-3 5e-3 -- 5e-3 5e-3 --
VC-1 sr 38.67 10.67 24.67 84.19 59.90 72.05
VC-1+IB sr 37.33 14.67 26.00 84.69 62.91 73.80
$\beta$ 1e-3 1e-3 -- 5e-3 5e-3 --
MPI sr 60.00 9.33 34.67 79.69 44.13 61.91
MPI+IB sr 61.33 12.00 36.67 79.91 46.78 63.34
$\beta$ 1e-4 1e-4 -- 5e-3 5e-3 --
Table 7: Task-wise Performance on DMControl of CortexBench. We evaluated 5 tasks using 3 random seeds and reported the average success rate (sr) and Lagrange multiplier $\beta$. The best performance is highlighted in bold.
@@
ResNet sr 38.32 88.37 92.20 91.42 64.34 74.93
ResNet+IB sr 50.75 90.42 99.78 96.39 87.37 84.94
$\beta$ 1e-3 1e-3 1e-3 1e-3 1e-3 --
ViT sr 7.22 3.39 18.97 18.05 4.43 10.41
ViT+IB sr 4.27 12.34 24.11 17.45 4.46 12.53
$\beta$ 1e-3 1e-3 1e-3 1e-3 1e-3 --
R3M sr 17.01 65.31 53.73 49.25 16.27 40.31
R3M+IB sr 16.19 72.10 57.34 46.48 16.60 41.74
$\beta$ 1e-3 1e-3 1e-3 1e-3 1e-3 --
Voltron sr 1.65 8.56 44.06 46.27 26.19 25.35
Voltron+IB sr 6.93 23.85 35.89 56.48 42.68 33.16
$\beta$ 1e-3 1e-3 1e-3 1e-3 1e-3 --
VC-1 sr 20.02 85.35 74.01 64.65 25.07 53.82
VC-1+IB sr 21.52 80.91 74.80 67.10 30.34 54.93
$\beta$ 1e-3 1e-3 1e-3 1e-3 1e-3 --
MPI sr 38.76 88.43 75.87 68.92 25.27 59.45
MPI+IB sr 33.82 86.44 86.99 69.29 30.50 61.41
$\beta$ 1e-3 1e-3 1e-3 1e-3 1e-3 --
Table 8: Task-wise Performance on Meta-World of CortexBench. We evaluated 5 tasks using 3 random seeds and reported the average success rate (sr) and Lagrange multiplier $\beta$. The best performance is highlighted in bold.
@@
ResNet sr 40.00 74.67 94.67 100.00 96.00 81.07
ResNet+IB sr 49.33 76.00 94.67 100.00 96.00 83.20
$\beta$ 1e-4 1e-3 1e-3 1e-3 1e-4 --
ViT sr 13.33 13.33 21.33 37.33 73.33 31.73
ViT+IB sr 13.33 9.33 18.67 62.67 76.00 36.00
$\beta$ 1e-4 1e-4 1e-4 1e-4 1e-4 --
R3M sr 42.67 56.00 38.67 66.67 61.33 53.07
R3M+IB sr 38.67 53.33 38.67 68.00 72.00 54.13
$\beta$ 1e-3 1e-3 1e-4 1e-4 1e-3 --
Voltron sr 60.00 58.67 68.00 82.67 93.33 72.53
Voltron+IB sr 57.33 74.67 54.67 93.33 92.00 74.40
$\beta$ 1e-4 1e-4 1e-4 1e-4 1e-4 --
VC-1 sr 68.00 60.00 65.33 100.00 94.67 77.60
VC-1+IB sr 70.67 76.00 69.33 100.00 96.00 82.40
$\beta$ 5e-3 5e-3 5e-3 5e-3 5e-3 --
MPI sr 61.33 40.00 58.67 100.00 72.00 66.40
MPI+IB sr 61.33 53.33 58.67 100.00 73.33 69.33
$\beta$ 1e-3 1e-3 1e-3 1e-3 1e-3 --
Table 9: Performance on language-condition multi-task benchmark LIBERO. We evaluated 40 tasks of 4 suites using 3 random seeds and reported the average success rate (sr) and Lagrange multiplier $\beta$. The best performance is bolded.
@@ C.1 Details of Simulation Experiments
We provide the task-wise results and corresponding $\beta$ values for CortexBench. The results for Adroit and TriFinger in Table 6, DMControl in Table 7, and MetaWorld are shown in Table 8. Across almost all tasks in CortexBench, incorporating the IB consistently improves performance compared to vanilla BC methods. Notably, models such as ResNet+IB, VC-1+IB, and MPI+IB often achieve the highest success rates, demonstrating the benefits of redundancy reduction in latent representations. In most cases, properly tuning $\beta$ (e.g., selecting values in the range of 1e-4 to 5e-3) leads to noticeable improvements.
We also provide the corresponding $\beta$ values for LIBERO in Table 9. Across all suites in LIBERO, incorporating the IB consistently improves performance over vanilla BC methods. The chosen values of $\beta$ (e.g., 1e-4 to 5e-3) effectively balance compression and predictive power.

### v4

@@ 1 Introduction
Behavior Cloning (BC), one of the simplest and most widely used methods in Imitation Learning (IL), learns a mapping from states to actions by training on state-action pairs from expert demonstrations. BC has been widely studied in autonomous driving (Bain & Sammut, 1995; Torabi et al., 2018), robotics control (Argall et al., 2009) and game AI (Pearce & Zhu, 2022). In robot manipulation, BC has become a foundational approach, enabling robots to replicate expert actions based on sensory inputs such as images or proprioception information like gripper states. To enhance the generalization of robots, most BC methods focus on incorporating large datasets of human or manipulation videos (Jang et al., 2022; Zhang et al., 2024; Cheang et al., 2024; Saxena et al., 2025; Cui et al., 2025), or integrating additional text and visual information (Jia et al., 2024; Wen et al., 2024a; Hu et al., 2024; Zhang et al., 2025b). While these methods have made significant progress in improving generalization by leveraging more diverse information, they often neglect a critical aspect: whether the learned representations contain significant redundant information.
@@ 2.1 Behavior Cloning in Robot Manipulation
Behavior Cloning (BC), first introduced by (Pomerleau, 1991), is a well-known Imitation Learning (IL) algorithm that learns a policy by directly minimizing the discrepancy between the agent’s actions and those of the expert in the demonstration data. To learn more generalizable representations, one class of visual representation learning methods pre-trains on large video datasets of robotics or humans, enabling rapid application of the pre-trained encoder to downstream robotic tasks. Notable examples include VC-1 (Majumdar et al., 2023), R3M (Nair et al., 2023), and Voltron (Karamcheti et al., 2023) . Meanwhile, another line of research focuses on training on even more extensive and diverse datasets with larger models, such as Internet-scale visual question answering and robot trajectory data (Brohan et al., 2023), as well as a vast collection of Internet videos (Cheang et al., 2024). Additionally, some methods further enhance generalization by incorporating additional sources of information. These include inferring textual descriptions based on the robot’s current state (Zawalski et al., 2024), leveraging visual trajectories (Wen et al., 2024a) and generated images (Tian et al., 2025; Hu et al., 2025; Zhang et al., 2025a), and integrating 3D visual information (Goyal et al., 2023). However, these methods have not deeply analyzed the redundancy in learned latent representations, and most also lack a solid theoretical foundation. Thus we extend the Information Bottleneck (IB) principle to BC, addressing this fundamental gap.
@@ 4.1 Model Architecture
As illustrated in Figure 2 (b), spatial fusion involves extracting spatial features from data at a given time step or concatenating features across multiple time steps along the feature dimensions. This approach does not explicitly differentiate between time steps but instead processes the aggregated features as a whole, emphasizing the modeling of inter-feature relationships. The spatial fusion module can be implemented using Multi-Layer Perceptrons (MLPs), Convolutional Neural Networks (CNNs), Spatial Transformers, or even simple concatenation operations. These methods are primarily designed to learn highly generalizable visual encoders by leveraging large-scale human video datasets. The pretrained encoders are then fine-tuned for downstream robotic tasks (Majumdar et al., 2023; Zeng et al., 2024).
@@
The latent representation $Z$, which integrates both spatial and temporal information, is then passed through a policy head to generate actions. Policy heads typically include MLPs, GMMs, diffusion policy (Chi et al., 2023; Reuss et al., 2024), and other designs (Gong et al., 2024). We choose an MLP for clarity and ease of empirical analysis.
@@ 4.2 Behavior Cloning with Information Bottleneck
where $z_{t}=F(x_{t})$ and $F(\cdot)$ denotes the fusion module. When $\beta<0$, the information from $X$ to $Z$ increases; when $\beta=0$, the method reduces to vanilla BC; and when $\beta>0$, the information from $X$ to $Z$ decreases.
@@ 5.1 Embodied Evaluation
Implementation. In CortexBench, for four partial fine-tuning methods, we train for 100 epochs on each task using the Adam optimizer with a learning rate of 1e-3, a batch size of 512, and weight decay of 1e-4, with learning rate decay applied using a cosine annealing schedule. For two full fine-tuning methods, we train for 50 epochs with a learning rate of 1e-4 and a batch size of 256. In LIBERO, we train for 50 epochs using the AdamW optimizer with a learning rate of 1e-4 and a batch size of 64, decayed using a cosine annealing schedule. In real-world evaluation, we train VC-1 for 200 epochs using the Adam optimizer with a learning rate of 1e-3 and a batch size of 512. CogAct is trained for 8k steps with the AdamW optimizer, using a learning rate of 2e-5 and a batch size of 128. For BC+IB methods, the model used in MINE consists of a two-layer MLP, with a learning rate of 1e-5. The Lagrange multiplier in Equation 6 ranges from 1e-4 to 1e-2 in this work.
@@ 5.3 Performance on LIBERO
Finding 8: IB is particularly effective for tasks requiring diverse feature extraction, such as distinguishing distinct task objectives or differentiating between various objects, as in LIBERO-Goal and LIBERO-Object. By filtering out irrelevant information, IB facilitates better generalization and more compact representations. However, the impact is less pronounced in spatial tasks such as LIBERO-Spatial, which rely heavily on structural information that can be disrupted by excessive compression. For long-horizon tasks, the main performance bottleneck lies in the lightweight baseline models, whose limited capacity restricts their effectiveness on more complex tasks such as LIBERO-Long.
@@ 5.5 More Analysis
Effect of the Lagrange multiplier $\beta$ of Equation 6. This experiment evaluates how incorporating IB enhances performance. Since the MINE model’s parameters are fixed, the key difference between BC+IB and BC lies in the parameter $\beta$, which balances compression and predictive power. For LIBERO experiments, $\beta$ is explored within 1e-4, 1e-3, 5e-3, 1e-2. As shown in Figure 5, IB improves performance within a specific $\beta$ range, with a peak observed at an undetermined value. However, across all experiments, $\beta$ around 1e-4 consistently yields stable improvements.
@@ 6 Limitations and Discussion
While our work provides extensive experimental validation of the effectiveness of IB and the necessity of input redundancy reduction in robotics representation learning, several limitations remain. First, we do not comprehensively assess the scalability of the IB approach in large-scale robotic models. Most of our experiments are conducted using relatively lightweight architectures, and only the real-world experiments employ CogAct. We have not systematically evaluated more complex models such as vision-language-action (VLA) architectures (Brohan et al., 2023; Song et al., 2025; Zhao et al., 2025) due to the high computational and time requirements. Second, we do not explore alternative policy head designs, including transformer-based heads or VLA models (Kim et al., 2024; Ding et al., 2024) that remove the explicit policy module and instead treat actions as text tokens. These architectures are left for future investigation. Third, although our method is evaluated on various benchmarks, its robustness to domain shifts, including changes in environment and task configurations, remains insufficiently studied. We hope this work inspires future research and advances the development of these methods.
7 Conclusion
In this study, we investigated the redundancy in latent representations for behavior cloning in robot manipulation and introduced the Information Bottleneck (IB) principle to mitigate this issue. By incorporating IB, we aimed to filter out redundant information in latent representations while preserving task-relevant features. Extensive experiments across various visual representation learning methods on CortexBench and LIBERO benchmark revealed insightful findings and demonstrated that IB consistently improves performance across diverse tasks and architectures. We hope this work will inspire future research to further integrate information-theoretic principles into robotics, not only as an optimization tool, but also as a framework for theoretical understanding and model design.
@@ Acknowledgements
This work was supported by the National Natural Science Foundation of China (Grant No. U21A20485), the National Science and Technology Innovation 2030 – Major Project (Grant No. 2022ZD0208800), the NSFC General Program (Grant No. 62176215, 62436005), and the Fundamental Research Funds for Xi’an Jiaotong University (Grant No. xzy022024012).
@@
This paper presents work whose goal is to advance the field of Machine Learning in Robotics. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ B.2.2 Baselines in LEBERO
Lagrange multiplier $\beta$ [1e-4, 1e-2]
@@ B.3.2 LIBERO
For BC+IB, all training and model parameters remain identical to those of BC, except for the IB-specific parameters.
@@ B.4.2 LIBERO
Table 6: Task-wise Performance on Adroit and Trifinger of CortexBench. We evaluated 4 tasks of 2 benchmarks using 3 random seeds and reported the average success rate (sr). The best performance is bolded.
@@
ResNet 65.33 66.67 66.00 87.12 56.06 71.59
ResNet+IB 69.33 74.67 72.00 87.14 57.45 72.30
ViT 61.33 9.33 35.33 78.77 32.37 55.57
ViT+IB 64.00 10.67 37.33 77.83 34.04 55.93
R3M 45.33 5.33 25.33 74.29 45.45 59.87
R3M+IB 52.00 2.67 27.33 75.04 46.23 60.63
Voltron 32.00 5.33 18.67 86.37 62.04 74.21
Voltron+IB 38.67 4.00 21.33 86.62 63.61 75.12
VC-1 38.67 10.67 24.67 84.19 59.90 72.05
VC-1+IB 37.33 14.67 26.00 84.69 62.91 73.80
MPI 60.00 9.33 34.67 79.69 44.13 61.91
MPI+IB 61.33 12.00 36.67 79.91 46.78 63.34
Table 7: Task-wise Performance on DMControl of CortexBench. We evaluated 5 tasks using 3 random seeds and reported the average success rate (sr). The best performance is highlighted in bold.
@@
ResNet 38.32 88.37 92.20 91.42 64.34 74.93
ResNet+IB 50.75 90.42 99.78 96.39 87.37 84.94
ViT 7.22 3.39 18.97 18.05 4.43 10.41
ViT+IB 4.27 12.34 24.11 17.45 4.46 12.53
R3M 17.01 65.31 53.73 49.25 16.27 40.31
R3M+IB 16.19 72.10 57.34 46.48 16.60 41.74
Voltron 1.65 8.56 44.06 46.27 26.19 25.35
Voltron+IB 6.93 23.85 35.89 56.48 42.68 33.16
VC-1 20.02 85.35 74.01 64.65 25.07 53.82
VC-1+IB 21.52 80.91 74.80 67.10 30.34 54.93
MPI 38.76 88.43 75.87 68.92 25.27 59.45
MPI+IB 33.82 86.44 86.99 69.29 30.50 61.41
Table 8: Task-wise Performance on Meta-World of CortexBench. We evaluated 5 tasks using 3 random seeds and reported the average success rate (sr). The best performance is highlighted in bold.
@@
ResNet 40.00 74.67 94.67 100.00 96.00 81.07
ResNet+IB 49.33 76.00 94.67 100.00 96.00 83.20
ViT 13.33 13.33 21.33 37.33 73.33 31.73
ViT+IB 13.33 9.33 18.67 62.67 76.00 36.00
R3M 42.67 56.00 38.67 66.67 61.33 53.07
R3M+IB 38.67 53.33 38.67 68.00 72.00 54.13
Voltron 60.00 58.67 68.00 82.67 93.33 72.53
Voltron+IB 57.33 74.67 54.67 93.33 92.00 74.40
VC-1 68.00 60.00 65.33 100.00 94.67 77.60
VC-1+IB 70.67 76.00 69.33 100.00 96.00 82.40
MPI 61.33 40.00 58.67 100.00 72.00 66.40
MPI+IB 61.33 53.33 58.67 100.00 73.33 69.33
Table 9: Performance on language-condition multi-task benchmark LIBERO. We evaluated 40 tasks of 4 suites using 3 random seeds and reported the average success rate (sr). The best performance is bolded.
@@ C.1 Details of Simulation Experiments
We provide the task-wise results and corresponding $\beta$ values for CortexBench. The results for Adroit and TriFinger in Table 6, DMControl in Table 7, and MetaWorld are shown in Table 8. Across almost all tasks in CortexBench, incorporating the IB consistently improves performance compared to vanilla BC methods. Notably, models such as ResNet+IB, VC-1+IB, and MPI+IB often achieve the highest success rates, demonstrating the benefits of redundancy reduction in latent representations. In most cases, properly tuning $\beta$ (e.g., selecting values in the range of 1e-4 to 1e-2) leads to noticeable improvements.